Shift summary eve 7/26/99

Philippe Gouffon (pgouffon@fdei01)
Tue, 27 Jul 1999 00:04:06 -0500

Shift Summary for eve shift, 26/jul/99

On Shift: Peter Shanahan, Philippe Gouffon

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7 (cosmic)

Accounting of lost time: ComEd power saving + MI problems

========================================================================
Runs ended this shift: 14034

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Boring & tedious, no beam

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14034 15:54 start of run 14034 (pedestal, to disk only (Valeri))
14034 16:09 end of run 14034
14035 16:22 start of run 14035 (cosmic, not to disk nor to tape)
17:27 ch13: startup in progress. csah button pushed in
19:35 DN put a new CO2 bottle on line
20:50 MI PS problems, no estimate, next update in one hour
21:12 MI PS problems, no estimate, next update in two hours
nice pictures of burnt ps on the MCR web page...

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

1) do a muon run.
2) change NM4AN polarity to positive
3) Eliott requested NOT to BOOT NCD's